OpenStudy (anonymous):

Okay... here we go. a/b(2x-12)=c/d

OpenStudy (anonymous):

what are you looking to do/find?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

i got ad/2bc + 6=x

OpenStudy (anonymous):

How though?

hero (hero):

What do you need to solve for?

OpenStudy (anonymous):

ok so you're going to distribute the b, so that you get 2xb-12b. Then rewrite your equation with that in place of what you had before

OpenStudy (anonymous):

Then after that, you need to cross multiply

OpenStudy (anonymous):

When you do that, you should get AD=C(2Bx-12B). tell me if you get confused on that.

OpenStudy (anonymous):

okay hold on. and im solving for x.

OpenStudy (anonymous):

Wait where does the a go?

OpenStudy (anonymous):

the A gets multiplied by the D...then you multiply the C by (2Bx-12B) to get the answer i gave you above

OpenStudy (anonymous):

This is so confusing.

OpenStudy (anonymous):

lol, give me a minute and ill type it up for you

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

\[a \div b(2x-12)= c \div d\]

OpenStudy (anonymous):

so this is your original equation. draw this with the a in the numerator and the c in the numerator. Then multiply the A and the D. See how they're across from each other? thats cross multiplying. then multiply the C and the (2Bx-12B) and put that answer on the other side of the equal sign

OpenStudy (anonymous):

okay! I get it. thanks.

hero (hero):

That's not exactly right @rjeffries96

hero (hero):

The equation is this: \(\large\frac{a(2x-12)}{b} = \frac{c}{d}\)

hero (hero):

Here are the steps: 1. Multiply both sides by variable b 2. Divide both sides by variable a 3. Add 12 to both sides 4. Divide both sides by 2 5. variable x is isolated